Can someone help me.
I need to solve the system by addition or substitution:

3x - 4y = 8
6x - 2y = 10
Thanks,
Lisa

Answer by chitra(359)

Hello Lisa,

The given system of equations are :

3x - 4y = 8 ------------ (1)

6x - 2y = 10 ------------ (2)

To solve them by addition or subtraction, we equalise any one of the terms. That is either the x or the y.

So, we multiply equation (2) by 2 and subtract equation (1) from (2).

12x - 4y = 20
3x - 4y = 8
- + -
---------------
9x = 12

==> x = 4/3

Plugging this value in one of the above equation, we get the value of y.

3x - 4y = 8

3(4/3) - 4y = 8

4 - 4y = 8

4 - 8 = 4y

-4 = 4y

This implies y = -1

Hence, the answer.